OpenStudy (anonymous):

Solve for y: 2/3 y + 4 = 6 My answer is y = 3 Am I right? :)

OpenStudy (agent0smith):

Check your answer by plugging in y=3 into the equation... ie replace y with 3, and see if the equation is true. (2/3) y + 4 = 6

OpenStudy (anonymous):

It didn't work

OpenStudy (agent0smith):

Show your work...

OpenStudy (agent0smith):

Because you very likely made a mistake.

OpenStudy (agent0smith):

\[\Large \frac{ 2 }{ 3}*\frac{ 3 }{1 }+4 =???\]

OpenStudy (anonymous):

Oh 3 does work! Thanks! :)